From owner-hst-status@stsci.edu Thu Mar 19 07:31:12 1998 X-Sender: cwolf@mail.hst.nasa.gov To: george.albright@hq.nasa.gov, sadams@estif2.gsfc.nasa.gov, barnbaum@stsci.edu, preston.m.burch.1@gsfc.nasa.gov, jconwell@ems.jsc.nasa.gov, john.h.campbell.1@gsfc.nasa.gov, bcooke@sssp.hst.nasa.gov, carpenter@tma1.gsfc.nasa.gov, aferro@as.arizona.edu, jeffrey.m.hanley1@jsc.nasa.gov, hst-status@stsci.edu, chenning@pop500.gsfc.nasa.gov, gsfc.hst.general@newsfeed.gsfc.nasa.gov, j.keith.kalinowski.1@gsfc.nasa.gov, ken.ledbetter@hq.nasa.gov, j_lesko@swales.com, jlegg@stx.com, jleibee@hst.nasa.gov, dleckrone@hst.nasa.gov, ann.c.merwarth.1@gsfc.nasa.gov, jmccullo@ems.jsc.nasa.gov, gmorrow@hst.nasa.gov, mgw@resource-intl.com, mnolan@pop100.gsfc.nasa.gov, Jan.Owings@gsfc.nasa.gov, kpulkkinen@v2pop.hst.nasa.gov, gilbert.roth@hq.nasa.gov, rrasche@as.arizona.edu, arankin@pop400.gsfc.nasa.gov, steven.l.smith1@jsc.nasa.gov, stdesk@eso.org, rthompson@as.arizona.edu, edward.weiler@hq.nasa.gov, vweyers@pop400.gsfc.nasa.gov, howard.j.wood.1@gsfc.nasa.gov Subject: Daily Report #2083 03/19/98 HUBBLE SPACE TELESCOPE DAILY REPORT #2083 PERIOD COVERED: 0000Z (UTC) 03/18/98 - 0000Z (UTC) 03/19/98 Daily Status Report as of 078/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. As described in yesterday's report and in HSTAR 6399, the acquisition for this proposal defaulted to fine lock on FGS-3 only. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.2 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned, and no anomalies were reported. 1.3 Completed NIC/1/2/3 7836 (Infrared Radiation from Middle-Aged Pulsars) The near Infrared Camera (Cameras One, Two and Three) was used to make infrared observations of the middle-aged pulsar PSR-B0656. The spectral distributions indicate that both surface thermal emission and non-thermal magnetospheric processes are present. These components need to be resolved before reliable tests of theories of neutron-star cooling and non-thermal evolution can be made. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.5 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Three Sets of STIS/CCD 7911 (STIS Galactic Plane Pure Parallel Program) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude survey observations in parallel with the prime science observations. Short exposures with the 50CCD imaging mode will provide the wavelength zero-point calibration. These data will be placed immediately into the Hubble Data Archive. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Two Sets of WF/PC-2 7625 (WF/PC-2 Cycle 7 Earth Flats) The WF/PC-2 was used to make a series of Earth flat calibration images. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The observations were completed as planned, and no anomalies were reported. 1.8 Completed NIC/2 7330 (The Fueling of Active Nucleii: A NICMOS Snapshot Survey) The Near Infrared Camera (Camera Two) was used to probe the dark matter distribution of NGC-1241 far from the center by observing intervening absorption from the MgII doublet (lambda 2800) in the spectrum of one background QSO located in a geometrically optimal position. HST observations are required because this feature is in a wavelength range not accessible from the ground, and the QSO magnitudes and expected strength of absorption make IUE observations unfeasible. The proposal completed as expected with no anomalies. 1.9 Completed NIC/2 7327 (Elliptical Galaxy Evolution at High Redshift: IR Imaging of 60 Distant Clusters) The Near Infrared Camera (Camera Two) was used to make infrared observations high redshifted elliptical galaxies in galactic clusters CL1144-123. By measuring the structural properties of high redshift cluster elliptical galaxies, it is hoped to resolve the controversy regarding the age and formation history of bright cluster ellipticals. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Two Sets of NIC/2 7226 (Search for Massive Jupiters) The Near Infrared Camera (Camera Two) was used to make a corographic search for massive planets around nearby, young main sequence star HD102982. Because of the extreme youth of this object, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed STIS/CCD 7564 (High Redshift Galaxy Companions To Q1202-0725) The Space Telescope Imaging Spectrograph (CCD) was used to make slitless mode observations of the high redshifted galaxy companion BR1202-07 to quasar Q1202-0725. A wavelength calibration image was also taken. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 9 The acquisition scheduled at 078/053154Z failed. The attitude was good, and there were no stop, scan step, or search radius flags; however, the FGSs were unable to maintain fine lock. HSTAR 6401 was written. NICMOS proposal 7901, STIS proposal 7911, and WF/PC-2 proposal 7909 (all occurring during the period of tomorrow's report) were affected since the take data flag remained down. Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 16 Successful : 16 2.3 Operations Notes: The SSR correctable EDAC errors were cleared two times using ROP SSR-1. The ENGREP was reset using ROP NS-5. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH